# KEN KANGDA WREN1

### v4

@@ 2.2.2 Connection with Familiar Price Dynamics
The above has a drift by design relative to reference belief $\{{\pi}^{+}_{t}\}$ (of dynamics (A.3)):
@@
note ${\mu^{\pi,\pi}_{t}}\equiv 0\equiv{\mu^{\Pi,\Pi}_{t}}$, confirming the martingale property of inference and model-risk pricing. Further, as inference volatility $\sigma^{\Pi,l}_{t}$ (29) drives price-volatility ((26-27)), the price-of-model-risk (24) peaks as model-risk induced price-volatility peaks.
@@
$+$ $r^{Z\Delta}_{t}dt\big{(}\mathbf{1}_{+}(B)-\Pi^{+}_{j}(B)\big{)}$ (32)
$+$ $S^{\Delta}(t)(\sigma^{\Pi}_{t})^{2}\big{[}(\sigma^{l}_{t})^{2}\big{(}\mathbf{1}_{+}(B)-\Pi^{+}_{t}(B)\big{)}+\sigma^{l}_{t}dw^{\prime}_{t}\big{]}.$ (33)
Under trivial $\{Z_{t}\}$ (firmed-up value process), the above becomes $dS_{t}(B)\propto d\Pi^{+}_{t}(B)$, making asset-pricing a pure matter of model-risk inference, which further reduces to the standard inferential test of Peskir and Shirayev (2006) under constant $\{r^{D\Delta}_{t}\}$ (signal) and $\{\sigma^{D}_{t}\}$ (noise).
@@
Practical cases often see $\check{R}_{t}\ll r^{Z\Delta}_{t}\ll S^{\Delta}(t)$: the $B$-sure dynamic (31) dominated by model-drift (32) dominated by model-inference (33); dispersion among such assets is driven by model-risk inference. It often holds also that $\text{F}^{Z/D}_{t}:=(\sigma^{lZ}_{t})^{2}/(\sigma^{lD}_{t})^{2}\ll 1$, so $(\sigma^{l}_{t})^{2}\approx(\sigma^{lD}_{t})^{2}$: prices dynamics are driven by data other than those on tangible (firmed-up) values.
Averaged under model-risk belief $\{{\pi}^{+}_{t}\}$, dominant price-dynamic (33) has a classical form:
$\sigma_{t}\bigg{(}\frac{\mu_{t}}{\sigma_{t}}+dw^{\prime}_{t}\bigg{)},$ (34)
@@
Expansion in $(K-1)$, given Remark 9, makes the classical relationship more explicit:
@@
The proportionality factor characterises ex-ante pricing through an intuitive and direct metric of asset-pricing (Remark 9). How its level comes about is a matter of economic theory (e.g. risk-aversion). Regardless, our derivation links it firmly to ex-ante beliefs, making the above relevant to bias studies and measurements under model-uncertainty.
@@ 3.1 The Joint-Hypothesis Issue
Let the $\{\mathbf{F}_{t}\}$-adapted objective conditional probabilities of model-risk outcomes be $\{\mathbf{p}^{B}_{t}\}$. The ex-post drift of our asset, $\{rp_{t}\}:=\{\Sigma_{B}\mathbf{p}^{B}_{t}{Y^{B}_{t}}\}-\{S_{t}\}$, assuming $B$-indifferent $B$-sure risk-premia $\{\check{rp}_{t}\}$ (Section 2.2.2) for clarity, satisfies:
@@
The LHS terms are each known/observable. The RHS terms are entangled: no conclusion on risk-pricing $\{B\_RP_{t}\}$ can be made without one on $\{\mathbf{p}^{b}_{t}\}-\{{\pi}^{b}_{t}\}$, namely, bias. For our model-risk, any realised process under a realised $B$-outcome corresponds to one run of (31-33) as if $\mathbf{p}^{b}_{t}\equiv 1\text{ or }0$. Many ’like’ observations are needed for an estimate of (37). Yet, given ’one-off’ and heterogeneous sources of model-uncertainties, how can this be done?
3.2 Assets Facing Potential Model-Change
@@
When model-sure, our asset is a simple drift with a static noise. Upon some trigger, it may face an objective chance $\mathbf{p}^{(\cdot)}_{0}$ of entering a new state (e.g. from $B=+$ to $B=-$). Denote any new state by ’1’, any status quo, ’0’, and status quo model-sure pricing, $\{S^{\text{{\bf{0}}}}_{t}\}$.
Change-trigger may or may not follow a known law. Our interest lies in what ensues after its arrival. It amounts to treating model-risk as a compound process and ignoring the part governing how/when model-risk arises (considered known, valued and reflected in model-sure prices). The average interval between model crises, say $\lambda^{-1}$, is presumed long enough for our one-off methodology; further, consider $t\ll\lambda^{-1}$ and $t\ll T$ henceforth.
@@
With the economic nature $sign[1{\text{{\bf{0}}}}]:=sign[Y^{1}_{t}-Y^{\bf{0}}_{t}]$ of potential change explicit, ignoring $t$-dependence of change impact $S^{\Delta}(t)$ for $t\ll T$ and $r^{Z\Delta}_{t}\ll S^{\Delta}(t)$, canonical pricing reads ((25)):
$S_{t}=S^{\text{{\bf{0}}}}_{t}+sign[1{\text{{\bf{0}}}}]S^{\Delta}\cdot\Pi^{1}_{t},\text{ }t\ll T,$ (38)
@@
where the given RNE belief process $\{\Pi^{1}_{t}\}\sim\{\pi^{1}_{t}\}$ can be characterised by risk-pricing parameter $K:=(\frac{O_{f}[\pi^{1}_{0}]}{O_{f}[{\Pi}^{1}_{0}]})^{sign[1{\text{{\bf{0}}}}]}$ (Corollary 1); let the market be competitive, so $K\in(1,2)$ (Remark 9).
3.3 The Representation of Status Quo Bias
The objective conditional probability process $\{\mathbf{p}^{1}_{t}\}$ governing change-risk is determined by the objective but unknown unconditional law $\mathbf{p}^{1}_{0}$ of model-risk. Its relationship against the RNE $\{\Pi^{1}_{t}\}$ and reference belief process $\{\pi^{1}_{t}\}$ can be characterised through the following:
$\rho K^{sign[1{\text{{\bf{0}}}}]}\equiv\frac{O_{f}[\mathbf{p}^{1}_{0}]}{O_{f}[{\Pi}^{1}_{0}]},$ (39)
$\rho:=\frac{O_{f}[\mathbf{p}^{1}_{0}]}{O_{f}[{\pi}^{1}_{0}]};$ (40)
@@ 3.3 The Representation of Status Quo Bias
for instance, if a change-risk is unpriced, then $K=1$; or if it is classical, with its law $\mathbf{p}^{1}_{0}$ known and so without model-uncertainty, then $\rho=1$.
@@
Parameter $\rho$, born of model-uncertainty, is not part of any ex-ante economic risk-discounting. The joint-hypothesis issue takes a concrete form here: observable prices reflect the product (39), and to split it into ’intention’ $K$ and ’mistake’ $\rho$ external criteria seem essential. This, as will be seen, need not be so; indeed, a high apparent bias under model-uncertainty may not be irrational either. The case of interest for us is when change is rare, $\mathbf{p}^{1}_{0}<\frac{1}{2}$, but priced as even rarer, such that (39-40) are both sizeable ($\rho\gg 1$). This ’justified strong status quo’ serves as a clear base against which ’change’ is defined and assessed (’unjustified strong status quo’, $\mathbf{p}^{1}_{0}>\frac{1}{2}$, allows easy wins to bets on ’change’ and so cannot persist).
3.4 Unconditional Model-Risk Symmetry, Dominance and Uniformity
Consider a market of a large number of assets subject to change-risks (e.g. new CEO, tort, war, recession). Let these risks be as characterised and economically priced in this market (e.g. due to their aggregate impact on market value/volatility/dispersion [^5] Dispersion is particularly intriguing in our context. As a priced factor it has been linked to market uncertainty, the business cycle and ’structural shifts’ (e.g. Demirer and Jategaonkar (2013), Kolari et al. (2021)).).
For clarity, let the unconditional all-cause probability of positive/negative economic change be even: at any time half of all potential changes in the market are positive/negative. Unconditional unevenness is addressed in Remark 17&18.
@@ 3.4 Unconditional Model-Risk Symmetry, Dominance and Uniformity
Let ’pricing intention’ $K$ ((39)) and ’apparent bias’ $\rho$ ((40)) be indifferent to the economic nature of potential change. Further, let parameter $K$, $\rho$ and change-impact $S^{\Delta}$ be uniform across the assets. Such uniformity amounts in practice to asset dispersions being driven by differences in inferential progression rather than in other aspects (Remark 11). Similarly, let all risk-inference be based on i.i.d data, with log-LR processes of fixed and uniform volatility. Note that each individual realisation of a change-outcome is considered random and independent still; while this may be a stretch for triggers such as war, it is a sensible start.
@@ 3.5 Conditional Cross-Sectional Trading
Pair-trading sorts assets into cohorts by cross-sectional statistics for long vs short positions. It amounts in our setup to conditioning by $B$-inferential events $\{\Pi^{1}_{t}=v\}$, $v\in(0,1)$, the driver of dispersion. By Bayes’ Rule (3) and definition (39), the $\{\Pi^{1}_{t}=v\}$-conditioned ex-post mean-excess $rp(\pm,v):=\pm(\mathbf{p}^{1}_{t}(v)-v)S^{\Delta}$ given the $sign[1{\text{{\bf{0}}}}]=\pm$ of potential change reads:
@@
where $(\sigma^{v})^{2}:=v\underline{v}$, making price-volatility the driver of excess ((29) and (34)). Note that the above is invariant under label-switching $1\leftrightarrow{\text{{\bf{0}}}}$ (i.e. $v\leftrightarrow\underline{v}$, +$\leftrightarrow$-+ ↔ - and $\rho\leftrightarrow\rho^{-1}$).
@@
With no uncertainty, $\rho=1$, the above is the ex-ante risk-premium when $\{\Pi^{1}_{t}=v\}$ ((24)): in classical markets, all the ’strategising’ merely earns what the market deems fair.
Despite excess (41) being driven by volatility, conditioning by $\{\Pi^{1}_{t}=v\}$, $v\in(0,1)$ under strong status quo amounts to ranking by momentum (past performance): $\{\Pi^{1}_{t}=v\}$ means $\{\Pi^{1}_{t}\}$ going from $\Pi^{1}_{0}\ll\frac{1}{2}$ to $v$ and so an excess-to-date of ±$\pm(v-\Pi^{1}_{0})S^{\Delta}$ ((38)). The associated model-uncertainty ($\propto(\sigma^{v})^{2}$) is a priori low and then evolves with dataflow. The window of opportunity for pair-trading is given by (C.8), Appendix C.2.
Drifts are hard to measure (e.g. Merton (1980)). Volatility trading is a good practical alternative. Volatility events, by (26-29), take the form $\{\Pi^{1}_{t}=v\}\cup\{\Pi^{1}_{t}=\underline{v}\}$, $v\in(0,\frac{1}{2}]$, and the $\{\sigma^{\Pi}_{t}=v\underline{v}\}$-conditioned ex-post mean-excess $rp_{t}^{\sigma}(v)$ averages over the $sign[1{\text{{\bf{0}}}}]$ of potential change and over the $\{\Pi^{1}_{t}=v\}$- and $\{\Pi^{1}_{t}=\underline{v}\}$-events in the volatility cohort:
@@
The relevant likelihoods (ratios) are determined by the underlying log-LR dynamics, whose cross-sections are Normal, making closed expressions possible. As a by-product, the ratio functions predict an association of volatility with ’negative momentum’ (Remark 17), a well-known phenomenon (e.g. Ang et al. (2006) and Wang and Xu (2015)).
The risk-reward curves of momentum trading (41) and volatility trading (42) are both concave in $v$ and vanishing as $v\underline{v}\to 0$.

### v6

@@ 2.2.2 Connection with Familiar Price-Dynamics
The above has a drift by design relative to reference belief $\{{\pi}^{+}_{t}\}$ (with dynamics (A.3)):
@@
note ${\mu^{\pi,\pi}_{t}}\equiv 0\equiv{\mu^{\Pi,\Pi}_{t}}$, reflecting the martingale property of inference and viable pricing. Further, as inference volatility $\sigma^{\Pi,l}_{t}$ (28) drives price-volatility (see (25-26)), price-of-model-risk $\{k^{\Pi}_{t}\}$ (by (23)) peaks where price-volatility peaks, ceteris paribus.
@@
$+$ $\big{[}\big{(}\mathbf{1}_{+}(B)-\Pi^{+}_{j}(B)\big{)}\cdot r^{Z\Delta}_{t}dt\big{]}$ (31)
$+$ $S^{\Delta}(t)(\sigma^{\Pi}_{t})^{2}\big{[}\big{(}\mathbf{1}_{+}(B)-\Pi^{+}_{t}(B)\big{)}\cdot(\sigma^{l}_{t})^{2}dt+\sigma^{l}_{t}dw_{t}\big{]}.$ (32)
@@
Practical cases often see $B$-sure dynamics (30) and model-drift (31) dominated by model-inference (32): $\check{R}_{t}\ll S^{\Delta}(t)$ and $r^{Z\Delta}_{t}\ll S^{\Delta}(t)$; indeed firmed-up values $\{Z_{t}\}$ are often nil or fixed so that asset-pricing is just a matter of model-risk inference, with $\{dS_{t}\}$ reduced to the inferential test equation of Peskir and Shirayev (2006) (for constant $\{r^{D\Delta}_{t}\}$ and $\{\sigma^{D}_{t}\}$). Further, it often holds that relative variance $\text{F}^{Z/D}_{t}:=(\sigma^{lZ}_{t})^{2}/(\sigma^{lD}_{t})^{2}\ll 1$ and so $(\sigma^{l}_{t})^{2}\approx(\sigma^{lD}_{t})^{2}$: model-risk is the prime price-mover, not tangible value. Dispersion among such assets is thus driven mainly by their differing underlying model-risk inference processes.
Averaged under model-risk belief $\{{\pi}^{+}_{t}\}$, dominant price-dynamic (32) has this familiar form:
$\sigma_{t}\bigg{(}\frac{\mu_{t}}{\sigma_{t}}dt+dw_{t}\bigg{)},$ (33)
@@
Expansion in (K1)(K-1), given Remark 7, recovers the standard and ubiquitous price-of-risk vs price-volatility relationship:
@@
The proportionality factor characterises ex-ante pricing through an intuitive and direct metric of asset-pricing (Remark 7). How its level comes about is a matter of economic theory (e.g. risk-aversion). Regardless, our derivation links it firmly to ex-ante beliefs, making the above relevant to bias studies and measurement under model-uncertainty.
@@ 3.1 The Joint-Hypothesis Issue
To focus on main issues, as in Section 2.2.2, consider usual-case assets under model-risk $B$ with $B$-independent $B$-sure risk-premia $\{\check{rp}(t)\}$. Let $\{\mathbf{p}^{B}_{t}\}$ be the $\{\mathbf{F}_{t}\}$-adapted true conditional probabilities of model-risk outcomes, given the true and known $B$-sure laws $\mathbf{W}^{T}_{B}$ and the true but potentially unknown unconditional law $\mathbf{p}^{B}_{0}$ of model-risk $B$. Its ex-post price-drift to be realised on average $\{rp_{t}\}:=\{\Sigma_{B}\mathbf{p}^{B}_{t}{Y^{B}_{t}}\}-\{S_{t}\}$ satisfies:
@@
The LHS terms are each observable or known, but the RHS are joint: no study of risk-pricing $\{B\_RP_{t}\}$ can be done without one on bias $\{\mathbf{p}^{b}_{t}\}-\{{\pi}^{b}_{t}\}$.
The above, in standard model-risk study settings, with the ’asset’ set to ’equity index’ and ’bias’ to nil, expresses the usual attribution of ’excess historical equity-return’ (LHS) to ’model-risk premium’ (RHS). Our focus is bottom-up and when bias is not necessarily nil.
3.2 Bias and Risk-Pricing Parameters under Model-Uncertainty
@@
An usual-case asset, if model-sure, is classical (a drift with a static noise). Upon the triggering of model-risk, it may enter a new state (e.g. from $B=+$ to $B=$-) with true probability $\mathbf{p}^{(\cdot)}_{0}$: denote the new state by ’1’, status quo, ’0’, and status quo model-sure pricing, $\{S^{\text{{\bf{0}}}}_{t}\}$.
The triggering events may or may not follow known laws. The process of how/when they arise is ignored, assuming that the average time between model crises, say some $\lambda^{-1}\gg 1$, is so long that each can be considered one-off. The period of interest is right after the arrival of a trigger but long before the next, so $t\ll\lambda^{-1}$, and long before horizon/maturity, so $t\ll T$.
@@
Canonical asset-pricing (24), with the direction $sign[1{\text{{\bf{0}}}}]:=sign[Y^{1}_{t}-Y^{\bf{0}}_{t}]$ of potential change explicit and the $t$-dependence of impact $S^{\Delta}(t)$ dropped given $t\ll T$ (and so $S^{\Delta}(t)\gg r^{Z\Delta}_{t}$), reads:
$S_{t}=S^{\text{{\bf{0}}}}_{t}+sign[1{\text{{\bf{0}}}}]S^{\Delta}\cdot\Pi^{1}_{t},$ (37)
@@
where the RNE belief process $\{\Pi^{1}_{t}\}\sim\{\pi^{1}_{t}\}$ is characterised by parameter $K:=(\frac{O_{f}[\pi^{1}_{0}]}{O_{f}[{\Pi}^{1}_{0}]})^{sign[1{\text{{\bf{0}}}}]}$ (Corollary 1); let the risk-pricing be competitive, so $K\in(1,2)$ (Remark 7).
The relationship between the true conditional probabilities $\{\mathbf{p}^{1}_{t}\}$ (given the true unconditional law $\mathbf{p}^{1}_{0}$), reference beliefs $\{\pi^{1}_{t}\}$ and RNE beliefs $\{\Pi^{1}_{t}\}$, all $\{\mathbf{F}_{t}\}$-adapted, can be characterised by:
$\rho:=\frac{O_{f}[\mathbf{p}^{1}_{0}]}{O_{f}[{\pi}^{1}_{0}]},\text{ and}$ (38)
$\rho K^{sign[1{\text{{\bf{0}}}}]}\equiv\frac{O_{f}[\mathbf{p}^{1}_{0}]}{O_{f}[{\Pi}^{1}_{0}]}.$ (39)
@@
If the change-risk is diversifiable, so unpriced, then $K=1$; if it is classical, so its law $\mathbf{p}^{1}_{0}$ is known, then $\rho=1$. Parameter $\rho$ stems from model-uncertainty and is thus unknown ex-ante.
@@
The joint-hypothesis problem takes a concrete form: observed prices reflect product (39), so external criteria seem necessary to separate ’intention’ $K$ from ’mistake’ $\rho$. We show that this need not be so. The case of interest for us is when change is unlikely, $\mathbf{p}^{1}_{0}<\frac{1}{2}$, but is priced as rare so that (39-38) are both sizeable ($\rho\gg 1$). Such a strong status quo forms a clear base against which ’change’ is defined and assessed. Unjustified strong status quo (i.e. $\mathbf{p}^{1}_{0}\geq\frac{1}{2}$ and $\rho\gg 1$) would allow easy wins for ’bet on change’ and so should not persist.
3.3 An Ideal Model-Uncertainty Driven Market
Consider a model-uncertainty driven market, of a large number of assets, each facing some change-risk $B$, whose nature may vary from asset to asset (e.g. new CEO, tort, war) but whose true $B$-sure laws are known. Let the $B$-risks be priced (i.e. correlated with the market’s stochastic-discount-factor), for their effects on aggregate value/volatility/cross-sectional dispersion [^6] Dispersion within members of an equity index, as a priced factor, has been linked to market uncertainty, the business cycle and structural shifts (e.g. Demirer and Jategaonkar (2013), Kolari et al. (2021)).. Further, let them be economically priced, so that by Proposition 1 their pricing is canonical with respect to some reference inferential belief process of the relevant $B$-risk.
Finally, let the realisation and detection of individual change-outcomes be stochastic and independent (even for macroeconomic triggers such as war or recession). Note that such a market, when seen as a whole, is an asset with non-binary model-uncertainty; this is so even when there is only one binary macroeconomic model-risk trigger.
@@ 3.3 An Ideal Model-Uncertainty Driven Market
For brevity, let the unconditional all-cause probability of positive/negative economic change be symmetrical: at any time half of all potential changes, of any origin, are positive/negative (unconditional asymmetry is addressed in Remark 14&15). Let ’pricing intention’ $K$ ((39)) and ’apparent bias’ $\rho$ ((38)) be indifferent to the economic direction and origin of change-risk, and the size of $K$, $\rho$ and risk-impact $S^{\Delta}$, uniform across the assets. Such uniformity in practice amounts to asset dispersions being driven by differences in the paths of model-risk inference rather than in any other aspect (Remark 9). Similarly, let all risk-inference be based on log-LR processes that are homogenous in time and uniform across.
@@ 3.4 Conditional Cross-Sectional Trading
Pair-trading sorts assets into cohorts by cross-sectional statistics for long vs short positions. It means in our setup conditioning by $B$-inferential events $\{\Pi^{1}_{t}=v\}$, $v\in(0,1)$, the driver of such statistics. Under (37-39), given the $sign[1{\text{{\bf{0}}}}]=\pm$ of potential change, the $\{\Pi^{1}_{t}=v\}$-conditioned ex-post mean-excess $rp(\pm,v):=\pm(\mathbf{p}^{1}_{t}(v)-v)S^{\Delta}$ unpacks (by Bayes’ Rule (3)) to:
@@
where $(\sigma^{v})^{2}:=v\underline{v}$, essentially making price-volatility (see (25-28)) the driver of excess. Note that the above is invariant under label-switching $1\leftrightarrow{\text{{\bf{0}}}}$ (i.e. $v\leftrightarrow\underline{v}$, $\leftrightarrow-$ and $\rho\leftrightarrow\rho^{-1}$).
@@
With no uncertainty, $\rho=1$, the above is by (23) the ex-ante risk-premium given $\{\Pi^{1}_{t}=v\}$: in classical markets all such strategising merely earns what the market deems fair.
Conditioning, or in practice, ranking, by events $\{\Pi^{1}_{t}=v\}$, $v\in(0,1)$, amounts to, under strong status quo, ranking by momentum (past performance): $\{\Pi^{1}_{t}=v\}$ means $\{\Pi^{1}_{t}\}$ going from $\Pi^{1}_{0}\ll\frac{1}{2}$ to $v$ and so an excess-to-date of $\pm(v-\Pi^{1}_{0})S^{\Delta}$ ((37)). The associated model-uncertainty ($\propto(\sigma^{v})^{2}$) is a priori low (’strong status quo’) and evolves with dataflow. For efficient pair-trading, the window of opportunity is given by (C.8), Appendix C.2.
Historical drifts are hard to measure in practice (e.g. Merton (1980)), so volatility trading can be an alternative. Volatility events $\{\sigma^{\Pi}_{t}=\cdot\}$ have the form $\{\Pi^{1}_{t}=v\}\cup\{\Pi^{1}_{t}=\underline{v}\}$, $v\in(0,\frac{1}{2}]$, by (25-28). Such $\{\sigma^{\Pi}_{t}=v\underline{v}\}$-conditioned ex-post mean-excess $rp_{t}^{\sigma}(v)$ is an average over the $sign[1{\text{{\bf{0}}}}]$ of potential change and over the $\{\Pi^{1}_{t}=v\}$- and $\{\Pi^{1}_{t}=\underline{v}\}$-events in the volatility cohort:
@@
The relevant likelihoods (ratios) are determined by the underlying log-LR dynamics, whose cross-sections are Normal, making closed expressions possible. As a by-product, these probability ratios predict an association of volatility with ’negative momentum’ (Remark 14), a well-known phenomenon (e.g. Ang et al. (2006) and Wang and Xu (2015)).
Observe that both the risk-reward of momentum (40) and volatility trading (41) are driven by $(\sigma^{v})^{2}$ essentially, and are thus concave as a function of $v$ and vanishing as $(\sigma^{v})^{2}\to 0$.